kekeman:

The aquarium has 3 more yellow fish than green fish. 60 percent of the fish are yellow. How many green fish are in the aquarium? Show your work.

QuestionCoveBot:

So say variable G represents green fish. So... \[\frac{ (G+3) }{ 60 }=\frac{ 40x }{ 120 }\] Cross multiply that.

QuestionCoveBot:

@kekeman

kekeman:

G = 20x - 3?

QuestionCoveBot:

No. (G+3) cross multiplies into 120, 60 into 40x (x is supposed to be g) It then becomes; \[60g=40g+120\] Subtract 40g on both sides.

QuestionCoveBot:

@kekeman

QuestionCoveBot:

Okay well after subtracting 40g on both sides, it becomes; \[20g=120\] Divide by 20; \[\frac{ 120 }{ 20 }\]Is what @kekeman ?

kekeman:

Which is 6

QuestionCoveBot:

Yes. So G = 6.

kekeman:

So there are 6 green fish in the aquarium

QuestionCoveBot:

Yes. And 9 yellow fish. So that makes 15 fish in total. Good job.

kekeman:

Thank you!

QuestionCoveBot:

You're welcome.